Grade 10Electromagnetic Induction

Resistance of 12 ohm is cut into three equal parts and add in triangular shape.find out resistance between two points of triangle

Ishaan
9 Years agoGrade 10

1 Answer

Vikas TU
9 Years ago
Resistance is directly proprtional to sq. of length.
Hence after cutting the length becomes l/3, l/3 and l/3 respectively.
Hence Resistances corresponding to those lengths would be:
R = kl^2
and 
(let) r = k(l/3)^2
R/r = 9kl^2/kl^2 => 9
r =  R/9 => 12/9 = > 4/3 ohm.
 
Now net  resistance between two points of triangle = >
8/3 and  4/3 in parallel.
=> (8/3) * (4/3) /((8/3) + (4/3))
=> (32/9) / 4
=> 8/9 ohm.